OpenStudy (anonymous):

X+Y=3 X-Y=5 solve each system. Write the answer as an ordered pair. please show steps!!!!

OpenStudy (anonymous):

add up

OpenStudy (anonymous):

or just think of two numbers that add to 5 and are three apart you will probably get it on the first try

OpenStudy (anonymous):

So do you solve independently of each other?

OpenStudy (anonymous):

think of two numbers that add up to 5

OpenStudy (anonymous):

for example, 2 and 3 add to 5 but \(3-2=1\) not \(3\) so that is not your answer try 1 and 4 \(4+1=5\) and \(4-1=3\) so that works

OpenStudy (mrnood):

add both sides of the equations together to make another equation That will allow you to solve x then use that value of x in either of the original equations - that will solve for y the solution is (x,y) (the ordered pair)

OpenStudy (anonymous):

or you can add in a column \[x+y=5\\ x-y=3\] add up and get \[2x=8\] making \(x=4\)

OpenStudy (anonymous):

x+y=3 if x =4 then (4, -1) x-y=5 if x =4 then (??